OpenStudy (anonymous):

Each edge of a cube is 5 meters long. Which is the volume of the cube? answers: 10 m^3 15 m^3 125 m^3 5 m^3

Parth (parthkohli):

You just cube that or raise it to the third power, as volume of a cube is Edge cubed.

OpenStudy (anonymous):

uh..

Parth (parthkohli):

What is \(5^3 = 5 * 5 * 5 = ? \)

OpenStudy (anonymous):

15

Parth (parthkohli):

Not 5 + 5 + 5, but 5 * 5 * 5.

OpenStudy (anonymous):

oh.. 125

Parth (parthkohli):

Yes!